OpenStudy (sissyedgar):

♥ Help Please! ♥ 1. Henrietta did an experiment. She started out with 800 bacteria cells. She found that the growth rate of the bacteria cells was 3.2%. Sketch the graph that represents the situation. Label the y-intercept and the point that represents the projected bacteria population 35 h from the time Henrietta started the experiment. ♥ Thank You! ♥

OpenStudy (sissyedgar):

@AloneS

OpenStudy (justin234):

what is the question? I only see info

OpenStudy (sissyedgar):

"Sketch the graph that represents the situation."

OpenStudy (mecharv):

I remember seeing a similar question in peeranswer 🤔

OpenStudy (tkhunny):

Information is insufficient or incomplete. A growth rate of "3.2%" doesn't mean anything. We need a time period. Hours? Minutes? Days?

OpenStudy (sissyedgar):

That there is all it gave me

OpenStudy (sissyedgar):

@3mar

OpenStudy (3mar):

Could you give me 3 min, please?

OpenStudy (sissyedgar):

Sure

OpenStudy (sissyedgar):

Ouuu, you're green now!!! Congratz! ♥

OpenStudy (3mar):

I am so Sorry for being late for you, sister! and thanks very much for your concern! Praise for Allah that I am green now!

OpenStudy (sissyedgar):

No problem

OpenStudy (tkhunny):

Well, if we ASSUME it mean "3.2% per hour", then we can proceed.

OpenStudy (3mar):

For such problem, you would deal with exponential growth or decay! are you familiar with it?

OpenStudy (3mar):

"She started out with 800 bacteria cells" that means: at the beginning of the experiment (t=0) the number of bacteria cells was 800 P=800 at t=0 "She found that the growth rate of the bacteria cells was 3.2%." that means: after on hour: bacteria cells numbers increased by (0.032*800)=25.6 then it became = 800+25.6=825.6 and so on! So you can represent that growth by the function: \[\Huge P=a*b^t\] where a=800 and t=1+0.032

OpenStudy (3mar):

@sissyedgar Did you get the idea? Try it! Put t=0, then t=1, .... what you got!?

OpenStudy (sissyedgar):

So sorry. I was in class!

OpenStudy (3mar):

No problem at all! I hope you did well in that class!

OpenStudy (3mar):

I am with you if you faced any difficulties! Let me know where you are back!

OpenStudy (sissyedgar):

Ok, I'm back. I ended up having back to back classes and worked on this problem a smidge with my teacher. We found the equation but I still have to graph it

OpenStudy (3mar):

\[\Huge {P=\color{DeepPink }a*\color{HotPink }b^t}\] \[\Huge {P=\color{DeepPink }{800}*\color{HotPink }{1.032}^t}\] and t in hours, and P is bacteria cells number. Try to graph it using this cool and interesting graphing calculator! If you still need help, I am at your service!

OpenStudy (sissyedgar):

We figured it out but thank you very much.. Though we found the equation to be....... p=800e^(3.2%*35)

OpenStudy (3mar):

Nice. Happy to know that. Any Help... Any Time...

OpenStudy (3mar):

I think it is just like that! http://openstudy.com/users/3mar#/updates/5832cc84e4b0080d89da0256